# hyperrefSuppressing link with empty target

### v1

@@ 3.5 Enthalpy Transport
where $\rho$ is the density, $T$ is the temperature, and $c_{p}$ is the specific heat. We plot the zonal, meridional, and vertical enthalpy flux for select pressure levels, chosen to emphasise differences in outer and deep atmosphere transport, for our three dynamical regimes in Figure 9.
@@
Starting in the slowly rotating regime, shown in the top row of Figure 9, we find that the zonal and meridional transport in the outer atmosphere (9(a)/c) is dominated by the divergent day-night wind, with eastward/westward enthalpy transport east/west of the sub-stellar point, and polewards/equatorwards meridional transport of the day/night side. Taken together, these flows suggest that enthalpy is transported from the irradiated day-side to slightly west of the anti-stellar point on the night-side, i.e. exactly where the night-side hot-spot eventually forms (5(d)).
The vertical advection of enthalpy in the outer atmosphere (9(d)) reinforces this link between horizontal transport and the night-side hot-spot, with a significant down-flow developing exactly where the night-side hot-spot forms, tracing the convergence of the horizontal winds. We also find a significant upwelling on the day-side, aligned with the hot-spot, leading to a global overturning-like circulation that is typical of atmospheres whose dynamics are shaped by divergent flows.
@@
The impact that the night-side hot-spot has on enthalpy transport becomes even more apparent at higher pressures. For example, in the mid-atmosphere, the zonal enthalpy transport (9(b)) is now no longer controlled just by the divergent day-night flows (although divergence driven dynamics remain dominant at high latitudes). Instead, we find significant, equatorial, enthalpy transport eastwards from the night-side towards the day-side/sub-stellar-point, driven by the ‘reversed’ day/night temperature difference. Note that this night to day transport essentially acts to close the overturning-like circulation described above, which has important implications for the deep atmosphere. Specifically, it means that the deep atmosphere is relatively quiescent and not as strongly linked to the outer atmosphere, leading to almost no deep heating and hence the suppression of advectively driven radius inflation (10(a)). Evidence for this can be seen in both 11(a) and 11(d), which show the vertical enthalpy transport against pressure for various longitudes/the global mean respectively: here we find a significant vertical enthalpy flux in the mid-atmosphere, around the same pressure as the night-side hot-spot, as well as a net (global mean) downward enthalpy flux in the deep atmosphere. Note that, despite the similarities in enthalpy flux magnitudes in local regions (compare 11(a) and 11(b)), we find that overall (i.e. when taking the global mean/calculating the net flux), the downward enthalpy advection is almost an order of magnitude lower here than in the classical hot Jupiter regime (compare 11(d) and 11(e)). This can be linked to the different vertical wind structures that the different horizontal wind profiles drive, see for example the much narrower downflow in 9(d) as opposed to 9(h).
@@
Finally, we note that the results of this enthalpy transport are clearly visible in the T-P profiles of the mid/outer atmosphere, as shown in 10(a). For example, on the night-side between $\sim 0.05\rightarrow\sim 2$\mathrm{bar}$$ we find clear evidence of the night-side hot-spot, including the significant cold-trap associated with the thermal inversion. We also find a reduced day-night temperature difference, suggesting that divergent day-night energy transport is highly efficient in the outer atmosphere, and that the day-night contrast of hot Jupiters might be sensitive to rotation rate.
@@
Moving onto the classical hot Jupiter regime, which we plot on the middle row of Figure 9, we find that the primary driver of enthalpy transport is now the rotational component of the wind. For example, at low pressures, we find that both the zonal and meridional enthalpy transport (9(e)/g) traces the eddy component of the wind (6(h)). As a result we find a significant asymmetry between westwards and eastwards day-night enthalpy transport: westwards of the sub-stellar point we find broad poleward transport which crosses the entire night-side, whereas eastwards of the sub-stellar point we find a much more longitudinally constrained profile linked with the off-equator wings that from in the outer and mid atmosphere (e.g. 5(e)). In turn, these off-equator wings are associated with a strong downward vertical enthalpy transport (9(h)) that is somewhat balanced by a broad upwelling near the sub-stellar point. Together forming a global circulation reminiscent of the global ‘overturning’ circulation.
@@
However unlike in the slowly rotating regime, we find that this global ‘overturning’ circulation extends deep into the atmosphere, where radiative forcing is weak, thus connecting the outer and deep atmospheres, and allowing for deep heating (10(b)). As for the reason why this circulation penetrates deeper into the atmosphere, this is because of the strong, deep zonal jet which drives strong mixing near the equator, matching what we find in the meridional circulation streamfunction (4(b)). Evidence for this deeper zonal jet can be seen in 9(f), which reveals strong eastwards advection near the equator at all longitudes, paired with weak off-equator transport linked to the equatorward pumping of eastwards angular momentum from mid-latitudes.
@@
Overall, these strong zonal/equatorial dynamics drive significant vertical enthalpy transport, as shown in 11(b)/e, where we find the highest net/global-mean downwards vertical enthalpy of our three regimes. A strong enthalpy flux which drives significant deep heating (see 10(b) and the dashed line in the deep atmosphere which shows the initial adiabat), suggesting that the planet is inflated relative to both more slowly rotating hot Jupiters and 1D atmospheric models (Sainsbury-Martinez et al. 2019).
@@
Finally we come to the rapidly rotating regime, which we plot on the middle row of Figure 9. Here we find that advective heat transport has become increasingly confined to low latitudes, with the equatorial jet acting as the main source of horizontal transport at almost all pressures. For example, in the outer atmosphere, 9(i)/k, we find horizontal enthalpy transport which is either driven by the jet (zonally) or by the very highly tilted and latitudinally constrained $m=1$ standing wave pattern (meridional). The jet transports enthalpy eastwards from the sub-substellar point, across the night-side, leading to the band of equatorial heating/mixing seen in 5(i), whilst the tilted standing wave primarily drives transport on the day-side (poleward/westward advection west of the sub-stellar point and equatorward/eastward advection to the east) and leads to the warped/compressed thermal butterfly also seen in 5(i).
@@
The influence of rotation on the enthalpy transport is also evident in the vertical enthalpy flow which is highly confined to low-latitude/equatorial regions. For example, as shown in 9(l) both the days-side upwelling and night-side downflow are confined to be within $\sim 20^{\circ}$ of the equator, which explains the lack of high-latitude downflows found at low Rossby numbers (7(b)). Further, due to the day-side confinement of the off-equator circulations, we find that the night-side downflows tend to be closer to the terminators/day-side than in the more slowly rotating regimes, although this changes as we move deeper into the atmosphere and the downflows become more tightly associated with a zonally-uniform zonal jet, as seen in 9(j).
@@
Overall, the Coriolis driven confinement of the horizontal and vertical enthalpy fluxes has a suppressing effect on the driving of deep atmospheric heating. That is to say that we find that, as shown in 11(c) downward enthalpy advection is significantly weaker than both of the other regimes considered here, suggesting that deep heating, and hence radius inflation, should be strongly suppressed. And the global-mean deep T-P profile confirms that this is indeed the case. However. if we confine our investigation of the deep atmosphere to equatorial regions, as we do in 10(c), we do find slight hints of deep heating. This occurs because the local vertical enthalpy advection is orders of magnitude stronger at the equator than elsewhere, and the latitudinal confinement for the off-equator circulations mean that latitudinal mixing is weak. With time, this mixing might lead to global heating of the deep atmosphere, albeit to a much lesser extent than in the classical regime, however a quick test with even weak radiative forcing back towards the initial adiabat (designed to reproduce the slight radiative cooling of the hot deep atmosphere) is more than enough to prevent this heating. As such, we feel confident in arguing that models in the rapid rotation regime should show little to no radius inflation, in agreement with the SDSS1411b results of Sainsbury-Martinez et al. (2021).
@@ 4 Non-Synchronous Rotation
Thus there is a strong possibility that next-generation exoplanet hunting missions will start to detect warm/hot Jupiters whose rotation period is not perfectly synchronised with their orbital period. This might have profound implications for their atmospheric dynamics. As such, we next explore how both small and large differences between the orbital and rotation period (i.e. non-synchronicity) affect the dynamics and resulting heat transport of such objects. Here our models fall into two regimes: the first is models in which the difference between the orbital ($\Omega_{\mathrm{orb}}$) and rotation ($\Omega_{\mathrm{rot}}$) angular rotation rates is a fraction of the overall rotation rate (e.g. $\Omega_{\mathrm{rot}}=0.9\Omega_{\mathrm{orb}}=0.9\Omega_{0}$ more generally when the difference is less than $\sim 15\%$ of the rotation rate - here that is the rotation rate of HD209458b), and the effect of non-synchronicity is relatively minor. And the second regime is when the difference is significant (e.g. $\Omega_{\mathrm{rot}}=4\Omega_{\mathrm{orb}}=4\Omega_{0}$), on the order of or greater than either the orbital or angular rotation rate, leading to significant disruption of the driving mechanisms discussed in section 3.
@@
Figure 12: Zonally and temporally averaged zonal wind profiles for four HD209458b-like models, two of which are non-synchronously rotating (top) and two of which are synchronously rotating (bottom), with rotation rates chosen to fall into our weak-non-synchronicity (left - $\Omega_{\mathrm{rot}}=0.9\Omega_{\mathrm{orb}}=0.9\Omega_{0}$ in the non-synchronous model and $\Omega_{\mathrm{rot}}=\Omega_{\mathrm{orb}}=0.9\Omega_{0}$ in its synchronous equivalent) and strong-non-synchronicity (right - $\Omega_{\mathrm{rot}}=4\Omega_{\mathrm{orb}}=4.0\Omega_{0}$ in the non-synchronous model and $\Omega_{\mathrm{rot}}=\Omega_{\mathrm{orb}}=4.0\Omega_{0}$ in its synchronous equivalent) regimes.
@@ 4.1 Zonal-Mean Zonal-Winds
We start in the weakly non-synchronous regime, where we find a zonal-mean zonal-wind (12(a)) that is very similar to its synchronously-rotating counterpart (12(c)), with differences on the order of that found when changing the averaging period of the temporal mean. This suggests that, when the difference between the rotational and orbital periods is small, the day-night temperature forcing that develops due to the strong stellar irradiation is sufficient to drive a long-lived standing Rossby and Kevin wave pattern leading to the development of a super-rotating jet equivalent to that found in the synchronous regime.
However as we increase the difference between the orbital and rotational angular rotation rates, we find that the differences between non-synchronous (12(b)) and synchronous (12(d)) models with the same planetary rotation rate starts to grow. In this highly non-synchronous regime, we find that: a) both eastwards (equatorial jet) and westwards (mid-latitude) winds are slower than in the synchronous regime and b) the weak eastwards equatorial jet that does develop is notably shallower than its synchronous counterpart, which in turn implies that vertical enthalpy advection/heating will be weakened/suppressed (see subsection 4.2 and subsection 4.4). These changes suggest that the time-dependent shift in the outer atmospheric irradiation (caused by the difference between the rotational and orbital period) hampers the formation of the standing wave pattern which drives both on and off-equator dynamics. We investigate the exact nature of these changes in subsection 4.3.
@@
Figure 14: Zonally and temporally averaged meridional circulation streamfunctions for four HD209458b-like models: two non-synchronous models (top), one in the weak-non-synchronicity regime (left) and one on the strong-non-synchronicity regime (right), paired with two synchronous models at the same rotation rates (bottom).

### v2

@@ 3.5 Enthalpy Transport
where $\rho$ is the density, $T$ is the temperature, and $c_{p}$ is the specific heat. We plot the zonal, meridional, and vertical enthalpy flux for select pressure levels, chosen to emphasise differences in outer and deep atmosphere transport, for our three dynamical regimes in Figure 7 and Figure 8.
@@
Starting in the slowly rotating regime, shown in the top row of Figure 7 and Figure 8, we find that the zonal and meridional transport in the outer atmosphere (7(a)/c) is dominated by the divergent day-night wind, with eastward/westward enthalpy transport east/west of the sub-stellar point, and polewards/equatorwards meridional transport of the day/night side. Taken together, these flows suggest that enthalpy is transported from the irradiated day-side to slightly west of the anti-stellar point on the night-side, i.e. exactly where the night-side hot-spot eventually forms (5(d)).
In order to better demonstrate the vertical advection of potential temperature/enthalpy in the outer atmosphere, and in particular the role that vertical winds play in forming the night-side hot-spot in the slowly rotating regime, we also plot longitudinal slices of the vertical wind, temperature, and vertical enthalpy flux in Figure 9. For example, we can see how vertical transport by a strong night-side downflow, and the associated adiabatic heating, leads to the formation of a hot-spot at around 0.2 to 0.5 bar in the slowly rotating regime. We can also see slight evidence for heating at pressures above the hot-spot as energy is deposited by the descending fluid parcels. Overall, we find a global overturning circulation, with a significant upwelling on the day-side, which is aligned with the upwelling, and the aforementioned downwelling on the night side. This is typical of atmospheres whose dynamics are shaped by divergent flows. Note that we find a similar overturning circulation in the classical hot Jupiter and rapidly rotating regimes, although here the advective heat transport is either deep (in the classical regime) or weak (in the rapidly rotating regime). This is in agreement with the differences in deep heating rates found between the three rotational regimes.
@@
The impact that the night-side hot-spot has on enthalpy transport becomes even more apparent at higher pressures. For example, in the mid-atmosphere, the zonal enthalpy transport (7(b)) is now no longer controlled just by the divergent day-night flows (although divergence driven dynamics remain dominant at high latitudes). Instead, we find significant, equatorial, enthalpy transport eastwards from the night-side towards the day-side/sub-stellar-point, driven by the ‘reversed’ day/night temperature difference. Note that this night to day transport essentially acts to close the overturning-like circulation described above, which has important implications for the deep atmosphere. Specifically, it means that the deep atmosphere is relatively quiescent and not as strongly linked to the outer atmosphere, leading to weak vertical enthalpy transport at high pressures (e.g. 8(c), where we find a vertical enthalpy flux an order of magnitude smaller than the flux found in the classical hot Jupiter regime; 8(f)) and hence the suppression of advectively driven radius inflation (10(a)). Evidence for this can be seen in both 11(a) and 11(d), which show the vertical enthalpy transport against pressure for various longitudes/the global mean respectively: here we find a significant vertical enthalpy flux in the mid-atmosphere, around the same pressure as the night-side hot-spot, as well as a net (global mean) downward enthalpy flux in the deep atmosphere. Note that, despite the similarities in enthalpy flux magnitudes in local regions (compare 11(a) and 11(b)), we find that overall (i.e. when taking the global mean/calculating the net flux), the downward enthalpy advection is almost an order of magnitude lower here than in the classical hot Jupiter regime (compare 11(d) and 11(e)). This can be linked to the different vertical wind structures that the different horizontal wind profiles drive, see for example the much narrower downflow in 8(a) as opposed to 8(d).
@@
Finally, we note that the results of this enthalpy transport are clearly visible in the T-P profiles of the mid/outer atmosphere, as shown in 10(a). For example, on the night-side between $\sim 0.05\rightarrow\sim 2$\mathrm{bar}$$ we find clear evidence of the night-side hot-spot, including the significant cold-trap associated with the thermal inversion. We also find a reduced day-night temperature difference, suggesting that divergent day-night energy transport is highly efficient in the outer atmosphere, and that the day-night contrast of hot Jupiters might be sensitive to rotation rate. This is in agreement with the analytic shallow-water models of Perez-Becker & Showman (2013), which showed that the day/night temperature contrast should increase with rotation rate (i.e. as the efficient overturning circulation is suppressed by rotational effects).
@@
Moving onto the classical hot Jupiter regime, which we plot on the middle row of Figure 7, Figure 8, and Figure 9, we find that the primary driver of enthalpy transport is now the rotational component of the wind. For example, at low pressures, we find that both the zonal and meridional enthalpy transport (7(d)/g) traces the eddy component of the wind (6(h)). As a result we find a significant asymmetry between westwards and eastwards day-night enthalpy transport: westwards of the sub-stellar point we find broad poleward transport which crosses the entire night-side, whereas eastwards of the sub-stellar point we find a much more longitudinally constrained profile linked with the off-equator wings that from in the outer and mid atmosphere (e.g. 5(e)). In turn, these off-equator wings are associated with a strong downward vertical enthalpy transport (8(d)) that is somewhat balanced by a broad upwelling near the sub-stellar point and to the east of the off-equator wings (8(e)). Together forming a global circulation reminiscent of the global ‘overturning’ circulation.
@@
However unlike in the slowly rotating regime, we find that this global ‘overturning’ circulation extends deep into the atmosphere, where radiative forcing is weak, thus connecting the outer and deep atmospheres, and allowing for deep heating (10(b) and 9(f)). As for the reason why this circulation penetrates deeper into the atmosphere, this is because of the strong, deep zonal jet which drives strong mixing near the equator, matching what we find in the meridional circulation streamfunction (4(b)). Evidence for this deeper zonal jet can be seen in 7(e), which reveals strong eastwards advection near the equator at all longitudes, paired with weak off-equator transport linked to the equatorward pumping of eastwards angular momentum from mid-latitudes.
@@
Overall, these strong zonal/equatorial dynamics drive significant vertical enthalpy transport, as shown in 8(f), 9(f), and 11(b)/e, where we find the highest net/global-mean downwards vertical enthalpy of our three regimes. A strong enthalpy flux which drives significant deep heating (see 10(b) and the dashed line in the deep atmosphere which shows the initial adiabat), suggesting that the planet is inflated relative to both more slowly rotating hot Jupiters and 1D atmospheric models (Sainsbury-Martinez et al. 2019).
@@
Finally we come to the rapidly rotating regime, which we plot on the bottom row of Figure 7, Figure 8, and Figure 9. Here we find that advective heat transport has become increasingly confined to low latitudes, with the equatorial jet acting as the main source of horizontal transport at almost all pressures. For example, in the outer atmosphere, 7(g)/k, we find horizontal enthalpy transport which is either driven by the jet (zonally) or by the very highly tilted and latitudinally constrained $m=1$ standing wave pattern (meridional). The jet transports enthalpy eastwards from the sub-substellar point, across the night-side, leading to the band of equatorial heating/mixing seen in 5(i), whilst the tilted standing wave primarily drives transport on the day-side (poleward/westward advection west of the sub-stellar point and equatorward/eastward advection to the east) and leads to the warped/compressed thermal butterfly also seen in 5(i).
@@
The influence of rotation on the enthalpy transport is also evident in the vertical enthalpy flow which is highly confined to low-latitude/equatorial regions. For example, as shown in 8(g) both the days-side upwelling and night-side downflow are confined to be within $\sim 20^{\circ}$ of the equator, which explains the lack of high-latitude downflows found at low Rossby numbers (17(b)). Further, due to the day-side confinement of the off-equator circulations, we find that the night-side downflows tend to be closer to the terminators/day-side than in the more slowly rotating regimes, although this changes as we move deeper into the atmosphere and the downflows become more tightly associated with a zonally-uniform zonal jet, as seen in 8(h)/i and 7(h). Evidence for this confinement of the downflows to the day-side can be seen in the left-hand column of Figure 9.
@@
Overall, the Coriolis driven confinement of the horizontal and vertical enthalpy fluxes has a suppressing effect on the driving of deep atmospheric heating. That is to say that we find that, as shown in 11(c) downward enthalpy advection is significantly weaker than both of the other regimes considered here, suggesting that deep heating, and hence radius inflation, should be strongly suppressed. And the global-mean deep T-P profile confirms that this is indeed the case. However. if we confine our investigation of the deep atmosphere to equatorial regions, as we do in 9(i) and 10(c), we do find slight hints of deep heating. This occurs because the local vertical enthalpy advection is orders of magnitude stronger near the equator than elsewhere (see 8(h)/i), and the latitudinal confinement for the off-equator circulations mean that latitudinal mixing is weak. With time, this mixing might lead to global heating of the deep atmosphere, albeit to a much lesser extent than in the classical regime, however a quick test with even weak radiative forcing back towards the initial adiabat (designed to reproduce the slight radiative cooling of the hot deep atmosphere) is more than enough to prevent this heating. As such, we feel confident in arguing that models in the rapid rotation regime should show little to no radius inflation, in agreement with the SDSS1411b results of Sainsbury-Martinez et al. (2021). Note that the slight northward shift of the downward enthalpy flux seen in 8(h)/i is driven by the slight asymmetry in the zonal winds, as seen in 1(c).
3.6 The Rossby Deformation Radius and Rhines Length Scales
Figure 12: A comparison of the maximum normalised Rossby deformation radius ($R_{rossby}$) and Rhines length ($R_{rhine}$) for eight of our synchronously rotating HD209458b-like atmospheric models which span the three rotation regimes discussed here. This figure can be compared with Figure 4 of Haqq-Misra et al. (2018).
We finish by briefly comparing our three rotational regimes to those proposed by Haqq-Misra et al. (2018) for terrestrial exoplanets. Haqq-Misra et al. (2018) use the non-dimensional Rossby Deformation Radius,
$R_{rossby}=\frac{NH}{2\Omega R_{p}}$ (18)
where H is the atmospheric scale height and N is the Brunt–Väisälä Frequency, and the Rhine Length,
$R_{rhine}=\frac{1}{R_{p}}\sqrt{\frac{U_{rms}R_{p}}{2\Omega}},$ (19)
to define three dynamical regimes. The slow rotation regime when both the normalised Rossby deformation radius ($R_{rossby}$) and Rhine length ($R_{rhine}$) are $>1$, the Rhine rotation regime when $R_{rossby}>1$ and $R_{rhine}<1$, and the rapidly rotating regime when both $R_{rossby}<1$ and $R_{rhine}<1$. They also propose a forth regime when $R_{rossby}<1$ and $R_{rhine}>1$, although none of their models fall into this regime. Each of these regimes are characterised by different dynamics, including global overturning circulation in the slowly rotating regime, significant equatorial super-rotation in the rapidly rotating regime, and a mix of dynamics in the Rhines regime. These regimes are broadly in line with the regimes we discuss here with a a global overturning circulation which dominates for slow rotators ($\Omega<0.5\Omega_{0}$) and a zonal-jet driven by the standing Rossby and Kelvin waves which dominates for rapid rotators ($\Omega\gtrsim 5\Omega_{0}$). However at intermediate rotation rates, our ‘classical’ hot Jupiter regime, instead of our models falling into the Rhines rotator regime, we instead fall into their theoretical fourth regime which, like the Rhines rotator regime, is characterised by a combination of a global overturning circulation and a standing-wave driven jet. This alignment with the regimes of Haqq-Misra et al. (2018) can be seen in Figure 12, where we plot the maximum values of the normalised Rossby deformation radius and Rhines length for eight of our synchronous models which span the three rotation regimes.
It is important to note that both the Rossby deformation radius and the Rhines length become very short in the rapidly rotating regime. In fact, depending upon the location in which the Rhines length is calculated, for our most rapidly rotating model ($\Omega=40\Omega_{0}$) the Rhines length can approach the resolution of our model: if we consider the global root mean square velocity ($U_{rms}$), we find that $R_{rhine}=2.5^{\circ}$, which is essentially identical to the horizontal extent of a grid cell in our models. However this value is so low in part because much of the atmosphere is quiescent in the rapidly rotating regime due to the suppression of off-equator motions by Coriolis forces. As such, if we instead constrain our calculation of $U_{rms}$ to winds within 10 degrees of the equator, we find that $R_{rhine}=6.0^{\circ}$, which we can resolve, if only just.
Consequently, given the interesting dynamics found here, we suggest that future studies investigate the rapidly rotating regime at a significantly higher resolution in order to fully capture all scales of the atmospheric dynamics at all latitudes.
@@ 4 Non-Synchronous Rotation
Thus there is a strong possibility that next-generation exoplanet hunting missions will start to detect warm/hot Jupiters whose rotation period is not perfectly synchronised with their orbital period. This might have profound implications for their atmospheric dynamics. As such, we next explore how both small and large differences between the orbital and rotation period (i.e. non-synchronicity) affect the dynamics and resulting heat transport of such objects. Here our models fall into two regimes: the first is models in which the difference between the orbital ($\Omega_{\mathrm{orb}}$) and rotation ($\Omega_{\mathrm{rot}}$) angular rotation rates is a fraction of the overall rotation rate (e.g. $\Omega_{\mathrm{rot}}=0.9\Omega_{\mathrm{orb}}=0.9\Omega_{0}$ more generally when the difference is less than $\sim 15\%$ of the rotation rate - here that is the rotation rate of HD209458b), and the effect of non-synchronicity is relatively minor. And the second regime is when the difference is significant (e.g. $\Omega_{\mathrm{rot}}=4\Omega_{\mathrm{orb}}=4\Omega_{0}$), on the order of or greater than either the orbital or angular rotation rate, leading to significant disruption of the driving mechanisms discussed in section 3. This work builds upon the non-synchronous models of, for example, Showman et al. (2009, 2015) by more thoroughly exploring the non-synchronous parameter space and investigating the effects of non-synchronous rotation on the vertical energy transport.
@@
Figure 13: Zonally and temporally, over five Earth-years of simulation time, averaged zonal wind profiles for four HD209458b-like models, two of which are non-synchronously rotating (top) and two of which are synchronously rotating (bottom), with rotation rates chosen to fall into our weak-non-synchronicity (left - $\Omega_{\mathrm{rot}}=0.9\Omega_{\mathrm{orb}}=0.9\Omega_{0}$ in the non-synchronous model and $\Omega_{\mathrm{rot}}=\Omega_{\mathrm{orb}}=0.9\Omega_{0}$ in its synchronous equivalent) and strong-non-synchronicity (right - $\Omega_{\mathrm{rot}}=4\Omega_{\mathrm{orb}}=4.0\Omega_{0}$ in the non-synchronous model and $\Omega_{\mathrm{rot}}=\Omega_{\mathrm{orb}}=4.0\Omega_{0}$ in its synchronous equivalent) regimes.
@@ 4.1 Zonal-Mean Zonal-Winds
We start in the weakly non-synchronous regime, where we find a zonal-mean zonal-wind (13(a)) that is very similar to its synchronously-rotating counterpart (13(c)), with differences on the order of that found when changing the averaging period of the temporal mean. This suggests that, when the difference between the rotational and orbital periods is small, the day-night temperature forcing that develops due to the strong stellar irradiation is sufficient to drive a long-lived standing Rossby and Kevin wave pattern leading to the development of a super-rotating jet equivalent to that found in the synchronous regime. This is consistent with the super-synchronous (1.5 times; see Figure 14) models of Showman et al. (2009).
However as we increase the difference between the orbital and rotational angular rotation rates, we find that the differences between non-synchronous (13(b)) and synchronous (13(d)) models with the same planetary rotation rate starts to grow. In this highly non-synchronous regime, we find that: a) both eastwards (equatorial jet) and westwards (mid-latitude) winds are slower than in the synchronous regime and b) the weak eastwards equatorial jet that does develop is notably shallower than its synchronous counterpart, which in turn implies that vertical enthalpy advection/heating will be weakened/suppressed (see subsection 4.2 and subsection 4.4). These changes suggest that the time-dependent shift in the outer atmospheric irradiation (caused by the difference between the rotational and orbital period) hampers the formation of the standing wave pattern which drives both on and off-equator dynamics. We investigate the exact nature of these changes in subsection 4.3. This shift is similar to that found by Showman et al. (2009), a slower, shallower, equatorial jet driven by a (then) theorised weakening in standing wave driven dynamics. Furthermore, like Showman et al. (2009), we also find that a high latitude and low pressure eastward flows form when the angular rotation rate significantly exceeds the orbital rotations rate. Although here our flows are slower, likely as a result of differences in the radiative treatment.
@@
Figure 15: Zonally and temporally, over five Earth-years of simulation time, averaged meridional circulation streamfunctions for four HD209458b-like models: two non-synchronous models (top), one in the weak-non-synchronicity regime (left) and one on the strong-non-synchronicity regime (right), paired with two synchronous models at the same rotation rates (bottom).